# Inferring entropy production in many-body systems using nonequilibrium MaxEnt

### v1

@@ Abstract
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. EP characterizes departure from thermodynamic equilibrium in microscopic physical systems, where it quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5]. There is great interest in quantifying EP in high-dimensional complex systems with many degrees of freedom and/or long non-Markovian memory, such as nonequilibrium disordered networks [6], biological active matter [7, 8], and brain networks [9, 10, 11, 12, 13], but this is computationally and statistically difficult. Standard methods require reconstruction of trajectory probability distributions, an intractable problem for systems with many spatiotemporal degrees of freedom.
In this paper, we propose a method to infer EP in high-dimensional stochastic systems. To do so, we introduce a variational principle that is the nonequilibrium analogue of the well-known Maximum Entropy Principle (MaxEnt) from equilibrium statistical physics. The dual form of this variational principle leads to a tractable, low-dimensional optimization problem. This solution provides an approximation of trajectory-level EP and a lower bound on the average EP, directly from trajectory samples and without explicit knowledge of trajectory probabilities. The dual principle provides a higher-order thermodynamic uncertainty relation (TUR) [14, 15], which we use to derive further simple bounds on EP.
@@
The system’s stochastic dynamics over sequences of $T+1$ discrete timesteps are described by a “forward” trajectory distribution $p(\vec{{\bm{x}}})$, where each trajectory $\vec{{\bm{x}}}=({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{T})$ is a sequence of $T+1$ states. In general, $T=1$ suffices for stationary (first-order) Markovian systems, while larger $T$ may describe non-Markovian and non-stationary systems. The system is also associated with a “reverse” trajectory distribution $\tilde{p}(\vec{{\bm{x}}})$. The trajectory EP of trajectory $\vec{{\bm{x}}}$ and the average EP across all trajectories are defined as [1]
@@
where $\langle\ldots\rangle_{p}=\sum_{\vec{{\bm{x}}}}\ldots p(\vec{{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler divergence. EP quantifies the violation of time-reversal symmetry. It vanishes in equilibrium when trajectory statistics are indistinguishable under the forward and reverse distributions.
@@
We propose a lower bound on EP based on a nonequilibrium generalization of the MaxEnt variational principle. We assume that we have access to $d$ trajectory observables, represented by a vector-valued function $\bm{g}(\vec{{\bm{x}}})\in\mathbb{R}^{d}$. Natural choices of observables include low-order interaction terms, such as 1-point ($x_{i,t}$) and 2-point ($x_{i,t}x_{j,t^{\prime}}$) fluctuating correlations. We do not require antisymmetric observables $\bm{g}(\vec{{\bm{x}}})=-\bm{g}(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$.
To quantify the EP captured by the temporal irreversibility of observables $\bm{g}$, we minimize the KL divergence to the reverse distribution, with the constraint that the observables $\bm{g}$ have the same expectations as under the forward distribution:
@@
The upper bound $0\leq{\Sigma}_{\bm{g}}\leq{\Sigma}$ follows because the forward distribution $p$ always satisfies the constraints. ${\Sigma}_{\bm{g}}$ vanishes only when $\bm{g}(\vec{{\bm{x}}})$ have the same expectations under forward and backward processes. As we will see, it achieves its maximum ${\Sigma}_{\bm{g}}={\Sigma}$ only when trajectory EP can be expressed as a linear combination of observables $\bm{g}(\vec{{\bm{x}}})$.
Eq. (2) provides a natural and interpretable bound on the EP. However, estimating EP directly from this optimization is not practical in high-dimensional systems, because there is no simple way to evaluate $\tilde{p}(\vec{{\bm{x}}})$, nor to optimize over the exponential (or even infinite) number of parameters that specify the distribution $q$. We solve both problems by exploiting a remarkable information-theoretic duality. As shown in the End Matter, the optimization problem (2) has the dual formulation:
${\Sigma}_{\bm{g}}=\max_{{\bm{\theta}}\in\mathbb{R}^{d}}\left({\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{p}-\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\right),$ (3)
It may be understood as an optimization over the Lagrangian multipliers ${\bm{\theta}}$ conjugate to constraints on expectations of $\bm{g}$.
Eq. (3) has two important features. First, it avoids estimation of trajectory probabilities, relying only on expectations of $\bm{g}(\vec{{\bm{x}}})$ under $p$ and $e^{{\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})}$ under $\tilde{p}$, which can be easily computed from empirical samples of observables $\bm{g}(\vec{{\bm{x}}})$.
In addition, Eq. (3) is a low-dimensional optimization problem, involving only as many parameters as observables in $\bm{g}$. For example, for a system with 1000 binary spins measured at two timepoints, Eq. (2) is a constrained optimization with more than $\sim 2^{1000}$ parameters required to specify $q$. For example, it $\bm{g}$ encodes 2-point correlations, then Eq. (3) is an unconstrained convex optimization over $\sim 1000^{2}$ parameters, which is quite feasible for modern computers using standard numerical techniques. In practice, we found that optimizing (3) using standard methods with early stopping on held-out test data led to good performance [16].
So far, we made no assumptions regarding multipartite structure (i.e., only a single variable can change at a time). However, as we show in End Matter, for multipartite dynamics and multipartite observables $\bm{g}(\vec{{\bm{x}}})$, Eq. (3) can be split into $N$ separate optimization problems, each one involving only a proportional fraction of the parameters and observations. This has important numerical advantages, since the computational cost of solving Eq. (3) scales superlinearly with size.
From the perspective of MaxEnt modeling, the distribution $q^{*}$ that optimizes (2) lies on a manifold of probability distributions forming exponential family (see End Matter):
@@
Using this, we can interpret the optimal ${\bm{\theta}}^{*}$ in Eq. (3) as the optimal parameters in the maximum likelihood (ML) problem ${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}D(p\|q_{{\bm{\theta}}})$, since $q_{{\bm{\theta}}^{*}}$ satisfies the constraints in (2) and achieves the optimum $q^{*}=q_{{\bm{\theta}}^{*}}$ (see End Matter). This interpretation also leads to the generalized Pythagorean theorem from information geometry [17], which decomposes EP into two nonnegative KL terms:
@@
The term ${\Sigma}^{\perp}_{\bm{g}}:={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP that cannot be captured by observables ${\bm{g}}$. It can be expressed as the optimal value of the ML problem, ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(p\|q_{\bm{\theta}})$. To shed light on this expression, it is useful to define a larger exponential family, $\omega_{f}(\vec{{\bm{x}}}):=\tilde{p}(\vec{{\bm{x}}})e^{f(\vec{{\bm{x}}})}/\langle e^{f}\rangle_{\tilde{p}}$, which contains the forward process as $p=\omega_{\sigma}$. We then write the ML problem as
@@
It is clear that ${\Sigma}^{\perp}_{\bm{g}}=0$ if and only if $\sigma(\vec{{\bm{x}}})={\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ for all $\vec{{\bm{x}}}$, meaning that the trajectory EP can be expressed as linear combination of the trajectory observables $\bm{g}$. More generally, the quantity ${\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ can be considered as the best possible approximation of $\sigma(\vec{{\bm{x}}})$ in the span of $\bm{g}$.
By considering a hierarchy of observables, our approach can define a hierarchical decomposition of EP, similar in spirit to information-geometric decompositions used in equilibrium MaxEnt [17, 18]. For example, if $\bm{g}_{[k]}(\vec{{\bm{x}}})$ specifies the set of $k$-point interactions, then we have the decomposition $0\leq{\Sigma}_{1}\leq{\Sigma}_{2}\leq\dots\leq{\Sigma}$. Here $\Sigma_{m}=D(q^{*}_{m}\|\tilde{p})$ is the contribution from all interactions up to order $m$, defined as the solution of Eq. (2) with expectation constraints on $\langle\bm{g}_{[k]}\rangle$ for $k=1,\dots,m$. Due to the Pythagorean theorem, the EP captured specifically by $k$-point interactions is itself a KL divergence, $\Sigma_{m}-\Sigma_{m-1}=D(q^{*}_{m}\|q^{*}_{m-1})$. In this way, we may derive an interpretable decomposition of how different interactions contribute to EP.
@@
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [19]. Consider any trajectory observable that can be expressed as a linear combination $o(\vec{{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o(\vec{{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o(\vec{{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the fluctuation-discounted expectation of all such $o(\vec{{\bm{x}}})$, and it can be understood as a higher-order TUR [14] that constrains the variance and the higher-order cumulants. This contrasts with most existing TURs, which quantify fluctuations only using the variance [20, 21, 22].
@@
To relate our approach to variance-based TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$. This gives the weaker bound
@@
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}.$ (7)
@@
It also gives the trajectory EP approximation, $\hat{\sigma}_{\bm{g}}(\vec{{\bm{x}}}):=\hat{{\bm{\theta}}}^{\top}\bm{g}(\vec{{\bm{x}}})-\ln\langle e^{{\hat{{\bm{\theta}}}}^{\top}\bm{g}}\rangle_{\tilde{p}}$ which satisfies $\langle\hat{\sigma}_{\bm{g}}\rangle_{p}=\widehat{{\Sigma}}_{\bm{g}}$. The advantage of these estimates is that they do not require optimization, only the solution of a linear system. Moreover, these estimates correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ (see Supplemental Material (SM) [16]). The bounds become equivalent (${\Sigma}_{\bm{g}}=\widehat{{\Sigma}}_{\bm{g}}$) when $\bm{g}$ has Gaussian statistics.
@@
The optimization-free TUR in (7) may be compared to a quadratic TUR. For stationary systems without odd variables, we find the inequality (see End Matter)
@@
where $\mathsf{K}_{(p+\tilde{p})/2}$ is the covariance matrix of $\bm{g}$ under the mixture distribution $(p+\tilde{p})/2$. For anti-symmetric observables, Eq. (8) may be considered the discrete-time version of the multidimensional TUR from Ref. [22].
@@
The bounds (7)-(8) may require one to solve linear systems involving very large covariance matrices. For multipartite systems, it is often possible to split these bounds into $N$ bounds that involve smaller matrices (see End Matter).
@@
Related work.— Information-theoretic TURs of the general form (3) were first proposed in [23] and were further explored in [24, 25, 15]. However, these works did not explicitly consider the problem of estimating EP in high-dimensional systems using low-dimensional observables. Several other approaches to EP inference have employed large-deviations theory [26], compression algorithms [27, 28, 29], and deep learning of probability flows [30].
@@
Our Pythagorean theorem (5) is related to previous decompositions of EP in interacting systems [31] and systems with nonconservative forces [32, 15]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [33, 34, 35, 36], sometimes called “maximum caliber”  [37]. However, our approach differs from earlier work on MaxEnt. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have some samples. This makes standard techniques and approximation methods for large-scale MaxEnt problems (such as mean-field and Bethe approximations  [38, 39, 6]) not directly applicable. Second, we care not only about the parameters ${\bm{\theta}}$, as is often the case in MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
@@
Lynn et al. [40, 41, 12] proposed a way to decompose and bound EP using a different information-theoretic optimization. Although originally focused on local EP in multipartite systems, it can be generalized to arbitrary observables $\bm{g}$ as
${\Sigma}_{\bm{g}}^{\textrm{L}}$ $=\min_{q}D(q\|\tilde{q})\;\>\text{subject to}\;\>\langle\bm{g}\rangle_{q}=\langle\bm{g}\rangle_{p},$ (9)
@@
It is usually restricted to stationary system without odd variables, so that $\tilde{q}(\vec{{\bm{x}}})=q(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$. Like our bound ${\Sigma}_{\bm{g}}$, ${\Sigma}_{\bm{g}}^{\textrm{L}}$ can be used to generate hierarchical decompositions of EP.
Although Eq. (9) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As a result, the resulting optimization problems are dramatically different. In particular, the optimal distribution $q^{*}$ in Eq. (9) is not in an exponential family and it does not have full support in general (see End Matter). Therefore, its dual problem is an optimization over an exponential number of variables, representing the exponential number of Lagrange multipliers required to enforce nonnegativity constraints $q(\vec{{\bm{x}}})\geq 0$ for all $\vec{{\bm{x}}}$. For this reason, at least in their standard form, neither the primal nor dual formulation of Eq. (9) are solvable for large systems, requiring the use of complex dimensionality-reduction techniques [12]. Future work may explore whether Eq. (9) can be solved using other types of optimization techniques.
@@
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our approach using two examples: a nonequilibrium asymmetric Ising model [42, 43, 6], and in vivo spike data from the Neuropixels Visual Behavior repository [44]. In both systems, we focus on a single time step and binary variables $x_{i,t}\in\{-1,+1\}$, $i\in\{1,\dots,N\}$ with time $t\in\{0,1\}$. We assume that the systems are in steady state, so that samples from $\tilde{p}$ are given by time-reversal of samples from the forward process $p$. For both examples, we optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [45]. We use a validation dataset for early stopping of the optimization, and report EP estimates computed on a held-out test dataset. This avoids overfitting, especially in far-from-equilibrium regimes where many reverse transitions are not sampled. Details of the data, optimization, and analysis are found in the SM [16] and our code repository [46].
@@
where we introduce the spin-flip operator, $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$, and the flip probability for spin $i$,
@@
where $\beta$ is an inverse temperature and $w_{ij}$ are (typically asymmetric) coupling parameters. An advantage of this model is that the steady-state EP has a simple closed-form expression in terms of expectation values, providing the ground truth for our EP estimators. Specifically, ${\Sigma}=\beta\sum_{i\neq j}w_{ij}\langle{(x_{i,1}-x_{i,0})x_{j,0}}\rangle_{\rm st}$, where $\langle\cdot\rangle_{\rm st}$ indicates expectations under the stationary process $p$.

### v2

@@ Abstract
The central quantity of interest in nonequilibrium thermodynamics is entropy production (EP) [1]. EP characterizes departure from thermodynamic equilibrium in microscopic physical systems, where it quantifies the dissipation of thermodynamic free energy. More generally, EP provides an information-theoretic measure of temporal irreversibility, including in meso- and macroscopic systems [2, 3, 4, 5]. There is great interest in quantifying EP in high-dimensional complex systems with many degrees of freedom and/or long non-Markovian memory, such as nonequilibrium disordered networks [6], biological active matter [7, 8], and neural systems [9, 10, 11, 12, 13], but this is computationally and statistically difficult. Standard methods require reconstruction of trajectory probability distributions, an intractable problem for systems with many spatiotemporal degrees of freedom.
In this paper, we propose a method to infer EP in high-dimensional stochastic systems. To do so, we introduce a variational principle that is the nonequilibrium analogue of the well-known Maximum Entropy Principle (MaxEnt) from equilibrium statistical physics. The dual form of this variational principle leads to a tractable optimization problem whose solution provides an approximation of trajectory-level EP and a lower bound on the average EP — directly from trajectory samples and without explicit knowledge of trajectory probabilities. The dual principle also provides a higher-order thermodynamic uncertainty relation (TUR) [14, 15], which we use to derive further simple bounds on EP.
@@
The system’s stochastic dynamics over sequences of $T+1$ discrete timesteps are described by a “forward” trajectory distribution $p(\vec{{\bm{x}}})$, where each trajectory $\vec{{\bm{x}}}=({\bm{x}}_{0},{\bm{x}}_{1},\dots,{\bm{x}}_{T})$ is a sequence of $T+1$ states. In general, $T=1$ suffices for stationary first-order-Markovian systems, while larger $T$ may describe non-Markovian and non-stationary systems. The system is also associated with a “reverse” trajectory distribution $\tilde{p}(\vec{{\bm{x}}})$. The trajectory EP of trajectory $\vec{{\bm{x}}}$ and the average EP across all trajectories are defined as [1]
@@
where $\langle\ldots\rangle_{p}=\sum_{\vec{{\bm{x}}}}\ldots p(\vec{{\bm{x}}})$ is the expectation over $p$ and $D(\cdot\|\cdot)$ is the Kullback-Leibler (KL) divergence. EP quantifies the violation of time-reversal symmetry, and it vanishes in equilibrium when trajectory statistics are indistinguishable under the forward and reverse distributions.
@@
We propose a lower bound on EP based on a nonequilibrium generalization of the MaxEnt variational principle. We assume that we have access to $d$ trajectory observables, represented by a vector-valued function $\bm{g}(\vec{{\bm{x}}})\in\mathbb{R}^{d}$. Natural choices of observables include low-order interaction terms, such as 1-point ($x_{i,t}$) and 2-point ($x_{i,t}x_{j,t^{\prime}}$) correlations. We do not require antisymmetric observables $\bm{g}(\vec{{\bm{x}}})=-\bm{g}(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$.
To quantify the EP captured by the temporal irreversibility of observables $\bm{g}$, we find the distribution that has the smallest KL divergence to the reverse process, subject to having the same observable expectations as the forward process:
@@
${\Sigma}_{\bm{g}}$ is always nonnegative, and it vanishes only when the expectation of $\bm{g}$ is the same under the forward and reverse processes. Furthermore, it is straightforward to show that ${\Sigma}_{\bm{g}}\leq D(p_{G}\|\tilde{p}_{G})\leq{\Sigma}$, where $p_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ and $\tilde{p}_{G}(\bm{g}^{\prime})=\sum_{\vec{{\bm{x}}}}\tilde{p}(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}$ indicate the distributions of $\bm{g}$ under the forward and backward processes (see Supplemental Material (SM) [16]). In general, estimating $D(p_{G}\|\tilde{p}_{G})$ is as difficult as estimating ${\Sigma}$, since it requires inference of the high-dimensional distributions $p_{G}$ and $\tilde{p}_{G}$.
On the other hand, by exploiting a remarkable information-theoretic duality, we may calculate ${\Sigma}_{\bm{g}}$ without explicitly inferring any probability distributions. Specifically, as shown in the End Matter, the optimization problem (2) has the dual formulation
${\Sigma}_{\bm{g}}=\max_{{\bm{\theta}}\in\mathbb{R}^{d}}\left({\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{p}-\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\right).$ (3)
Eq. (3) is a tractable convex optimization problem over the Lagrangian multipliers ${\bm{\theta}}$ that enforce the $d$ expectation constraints on $\bm{g}$. Importantly, the objective depends only on the expectations of $\bm{g}(\vec{{\bm{x}}})$ under $p$ and $e^{{\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})}$ under $\tilde{p}$, which can be easily computed from empirical samples of the observables.
As an example, consider a system with 1000 binary spins measured at two timepoints. A naive estimate of ${\Sigma}$ would require inference of $\sim 2^{1000}$ probabilities. Instead, with $\bm{g}$ encoding 2-point correlations, Eq. (3) results in a tractable unconstrained optimization problem over $\sim 1000^{2}$ variables.
So far, we made no assumptions regarding multipartite structure (i.e., only a single variable can change at a time). However, as we show in End Matter, for multipartite dynamics and multipartite observables, Eq. (3) can be split into $N$ separate optimization problems, each one involving only a proportional fraction of the variables and observations. This can dramatically reduce the amount of computation and memory required to numerically solve Eq. (3).
The distribution $q^{*}$ that optimizes (2) lies in the exponential family,
@@
We can interpret the optimal ${\bm{\theta}}^{*}$ in Eq. (3) as the optimal parameters in the maximum likelihood (ML) problem ${\bm{\theta}}^{*}=\operatorname*{arg\,min}_{{\bm{\theta}}}D(p\|q_{{\bm{\theta}}})$, since $q_{{\bm{\theta}}^{*}}=q^{*}$ satisfies the constraints in (2) and achieves the optimum objective value (see End Matter). This interpretation also leads to the information-geometric Pythagorean theorem [17], which decomposes EP into two nonnegative KL terms:
@@
The term ${\Sigma}^{\perp}_{\bm{g}}:={\Sigma}-{\Sigma}_{\bm{g}}=D(p\|q^{*})$ quantifies the EP not captured by ${\Sigma}_{\bm{g}}$. It can be expressed in terms of a ML problem, ${\Sigma}^{\perp}_{\bm{g}}=\min_{{\bm{\theta}}\in\mathbb{R}^{d}}D(p\|q_{\bm{\theta}})$. To interpret this expression, we define a larger exponential family $\omega_{f}(\vec{{\bm{x}}}):=\tilde{p}(\vec{{\bm{x}}})e^{f(\vec{{\bm{x}}})}/\langle e^{f}\rangle_{\tilde{p}}$ that contains the forward trajectory distribution $p(\vec{{\bm{x}}})=\omega_{\sigma}(\vec{{\bm{x}}})$. The ML problem can be written as
@@
From this expression, it is clear that ${\Sigma}^{\perp}_{\bm{g}}=0$ (and ${\Sigma}_{\bm{g}}={\Sigma}$) if and only if trajectory EP is a linear combination of the trajectory observables, i.e., $\sigma(\vec{{\bm{x}}})={\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ for all $\vec{{\bm{x}}}$. When ${\Sigma}^{\perp}_{\bm{g}}>0$, ${\bm{\theta}}^{*\top}\bm{g}(\vec{{\bm{x}}})$ provides the best possible approximation of $\sigma(\vec{{\bm{x}}})$ in the span of $\bm{g}$.
Eq. (4) implies that the conditional distribution of the optimal $q^{*}$ obeys $q^{*}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=\tilde{p}_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})$, with conditional probabilities given by Bayes’ rule, $p_{\vec{X}|G}(\vec{{\bm{x}}}|\bm{g}^{\prime})=p(\vec{{\bm{x}}})\delta_{\bm{g}^{\prime},\bm{g}(\vec{{\bm{x}}})}/p_{G}(\bm{g}^{\prime})$. Using the chain rule of KL divergence, we may further decompose ${\Sigma}^{\perp}_{\bm{g}}$ into
${\Sigma}^{\perp}_{\bm{g}}=D(p_{G}\|q^{*}_{G})+D(p_{\vec{X}|G}\|\tilde{p}_{\vec{X}|G})\,.$ (7)
The first term in Eq. (7) vanishes when $q^{*}_{G}=p_{G}$; it is the additional EP that could be inferred by constraining all higher-order statistics (not just the mean) in Eq. (2). The second term is the EP that cannot be inferred from any statistics of $\bm{g}$.
Finally, our method supports a hierarchical decomposition of EP, where contributions from interactions of increasing order (e.g., pairwise, triplets) are successively added, resulting in a structure $0\leq{\Sigma}_{1}\leq{\Sigma}_{2}\leq\dots\leq{\Sigma}$, with $\Sigma_{k}=D(q^{*}_{k}\|\tilde{p})$ capturing contributions from functions $\bm{g}(\vec{{\bm{x}}})$ containing interactions up to order $k$. This structure mirrors information-geometric decompositions in equilibrium MaxEnt [17, 18] and allows EP to be broken down into interpretable components, each satisfying $\Sigma_{k}-\Sigma_{k-1}=D(q^{*}_{k}\|q^{*}_{k-1})$.
@@
Thermodynamic uncertainty relations (TURs).— Eq. (3) can be interpreted as a TUR that relates EP and fluctuations of trajectory observables [19]. Consider any trajectory observable that can be expressed as a linear combination $o(\vec{{\bm{x}}})={\bm{\theta}}^{\top}\bm{g}(\vec{{\bm{x}}})$ for some ${\bm{\theta}}\in\mathbb{R}^{d}$. The first term in (3) is the expectation of $o(\vec{{\bm{x}}})$, while the second term is the cumulant generating function (CGF) of $o(\vec{{\bm{x}}})$ under $\tilde{p}$. Thus, ${\Sigma}_{\bm{g}}$ bounds the “fluctuation discounted” expectation of all such $o(\vec{{\bm{x}}})$, giving a higher-order TUR [14] that constrains all cumulants of trajectory observables. This contrasts with quadratic TURs, which constrain only the mean and variance [20, 21, 22].
@@
To relate our approach to quadratic TURs, we expand the CGF as $\ln\langle e^{{\bm{\theta}}^{\top}\bm{g}}\rangle_{\tilde{p}}\approx{\bm{\theta}}^{\top}\left\langle\bm{g}\right\rangle_{\tilde{p}}+{\bm{\theta}}^{\top}\mathsf{K}_{\tilde{p}}{\bm{\theta}}/2$, where $\mathsf{K}_{\tilde{p}}$ is the covariance matrix of $\bm{g}$ under $\tilde{p}$. Plugging into (3) gives the approximate optimizer $\hat{{\bm{\theta}}}$ as the solution to the linear system $\mathsf{K}_{\tilde{p}}\hat{{\bm{\theta}}}=\left\langle\bm{g}\right\rangle_{p}-\left\langle\bm{g}\right\rangle_{\tilde{p}}\equiv\left\langle\bm{g}\right\rangle_{p-\tilde{p}}$. This gives the weaker bound
@@
${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ $:=\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\left\langle\bm{g}\right\rangle_{p}-\ln\left\langle e^{\left\langle\bm{g}\right\rangle_{p-\tilde{p}}^{\top}\mathsf{K}_{\tilde{p}}^{-1}\bm{g}}\right\rangle_{\tilde{p}}.$ (8)
@@
$\widehat{{\Sigma}}_{\bm{g}}$ can be shown to correspond to a single Newton-Raphson step for maximizing (3) starting from ${\bm{\theta}}=\bm{0}$ [16]; it has the advantage of not requiring optimization, only the solution a linear system. The bound ${\Sigma}_{\bm{g}}\geq\widehat{{\Sigma}}_{\bm{g}}$ becomes tight when $\bm{g}$ has Gaussian statistics.
@@
We may also compare our results to a quadratic TUR. For stationary systems without odd variables, we find the inequality (see End Matter)
@@
where $\mathsf{K}_{(p+\tilde{p})/2}$ is the covariance matrix of $\bm{g}$ under the mixture distribution $(p+\tilde{p})/2$. For antisymmetric observables, Eq. (9) may be considered the discrete-time version of the multidimensional TUR from Ref. [22].
@@
The bounds (8)-(9) may require solving linear systems involving very large covariance matrices. For multipartite systems, it is often possible to split these bounds into $N$ bounds that involve smaller matrices (see End Matter).
@@
Related work.— Eq. (3) is closely related to the Donsker-Varadhan representation from large deviations theory [23], previously used to estimate information-theoretic quantities in machine learning [24] and thermodynamic inference [25, 26, 27]. However, these works did not consider the problem of thermodynamic inference in high-dimensional systems using multivariate observables. Other approaches to EP inference have employed compression algorithms [28, 29, 30] and deep learning of probability flows [31].
Our Pythagorean theorem (5) is related to previous decompositions of EP in interacting systems [32] and systems with nonconservative forces [33, 34, 15]. Moreover, the general form of Eq. (2) is a Maximum Entropy (MaxEnt) problem over trajectory distributions [35, 36, 37, 38], sometimes called “maximum caliber” (MaxCal) [39]. Related techniques have been used to infer models from which the entropy flow can be estimated, though without ensuring a lower bound [40].
@@
Our approach differs in several other ways from earlier work on MaxEnt and MaxCal. First, we minimize the KL divergence relative to an unknown prior distribution $\tilde{p}$, from which we typically only have some samples. This makes standard approximation methods for large-scale MaxEnt problems (such as mean-field and Bethe approximations [41, 42, 6]) not directly applicable. Second, we care not only about the parameters ${\bm{\theta}}$, as is typical of MaxEnt “inverse problems”, but also about bounding EP by the quantity ${\Sigma}_{\bm{g}}$. Both of these issues are resolved by the dual formulation (3).
@@
Lynn et al. [43, 44, 12] proposed a way to decompose and bound EP using a different information-theoretic optimization. Although originally focused on local EP in multipartite stationarity systems, it can be generalized to arbitrary observables $\bm{g}$ and non-multipartite systems as
${\Sigma}_{\bm{g}}^{\textrm{L}}$ $=\min_{q,\tilde{q}}D(q\|\tilde{q})\;\>\text{subject to}\;\>\langle\bm{g}\rangle_{q}=\langle\bm{g}\rangle_{p}.$ (10)
@@
Here we use assumption of stationarity and no odd variables, so that $\tilde{q}(\vec{{\bm{x}}})=q(\reflectbox{$\vec{\reflectbox{${\bm{x}}$}}$})$. Like our bound ${\Sigma}_{\bm{g}}$, ${\Sigma}_{\bm{g}}^{\textrm{L}}$ can be used to generate hierarchical decompositions of EP.
Although Eq. (10) resembles Eq. (2), it differs in that it simultaneously optimizes both arguments of the KL divergence, $q$ and $\tilde{q}$. As we discuss in the End Matter, the optimal distribution $q^{*}$ in Eq. (10) is not in an exponential family and in general it lacks full support. For this reason, Eq. (10) does not have a tractable dual expression (i.e., analogous to our dual (3)), and there is no straightforward way to scale ${\Sigma}_{\bm{g}}^{\textrm{L}}$ to large systems.
@@
Examples: Nonequilibrium spin model and Neuropixels dataset.— We illustrate our approach using two examples: a nonequilibrium kinetic Ising model [45, 46, 6], and in vivo spike data from the Neuropixels Visual Behavior repository [47]. In both systems, we focus on a single time step and binary variables $x_{i,t}\in\{-1,+1\}$, $i\in\{1,\dots,N\}$ with time $t\in\{0,1\}$, and assume stationarity.
To calculate ${\Sigma}_{\bm{g}}$, we optimize Eq. (3) using gradient ascent with Barzilai–Borwein step sizes [48]. We perform early-stopping using held-out validation data, which avoids overfitting even in the far-from-equilibrium regime where many reverse transitions are not sampled. Reported EP estimates are computed on held-out test data. Details of the data generation, optimization, and analysis are found in the SM [16] and our code repository [49].
@@
Here we introduce the spin-flip operator, $(\bm{x}^{[i]})_{i}=-x_{i}$ and $(\bm{x}^{[i]})_{j}=x_{j}$ for $j\neq i$. The flip probability for spin $i$ is
@@
where $\beta$ is an inverse temperature and $w_{ij}$ are (typically asymmetric) coupling parameters. We consider the diluted version of the model [50, 51, 52] with $k$ average neighbors. Here $w_{ij}=c_{ij}z_{ij}/\sqrt{k}$, with binary connections $c_{ij}\sim\textrm{Bernoulli}(k/(N-1))$ and real-valued weights $z_{ij}\sim\mathcal{N}(0,1)$.
The steady-state EP in this model has a simple closed-form expression, ${\Sigma}=\beta\sum_{i\neq j}w_{ij}\langle{(x_{i,1}-x_{i,0})x_{j,0}}\rangle_{\rm st}$, where $\langle\cdot\rangle_{\rm st}$ indicates expectations under the stationary process $p$. This provides ground truth to evaluate our estimators.